Louisiana State University Health Sciences Center Administration & Finance
 

Monday, May 22, 2006

  • The Roman Street Parking Garage emergency power panel, which controls the electricity to the 1st floor lighting and emergency lights in the stairwells, has been repaired and is re-energized.
  • The environmental remediation contractor completed the cleaning of the walkway by Stanislaus Hall. They are cleaning ducts on the 8th floor of 1542 Tulane Ave. and are continuing with the asbestos abatement in the basement. Cleaning has also begun on the 7th floor.
  • Building steam is now available to the autoclaves in both the CSRB and the MEB.
  • Upgrades to the MEB 3rd floor Seminar Rooms continue. The carpeting is almost complete. The installation of new lights is approximately 70% complete. Painting has been substantially completed; only touch up work remains to be done. One of the new white boards needs to be replaced. New ceiling tiles have been delivered and will be installed when the electrical work has been completed. The final improvement will be the installation of wireless network access points above the ceiling tiles.
  • The installation of new carpeting in Medical Center Stores has been completed. The furniture, shelving and stock are being moved back into the original space.
  • The installation of new carpeting in Campus Office Stores is approximately 50% complete.
  • Repairs to the emergency generators are progressing. The contractor installed the transfer switches at the Residence Hall over the weekend. A full crew will be on site to work on the generator at the MEB tomorrow.
  • The State Fire Marshal inspected the fire alarm (and related) systems at the Residence Hall today and granted the approval to reoccupy when the remainder of the preparations are completed in student rooms.
  • The State Fire Marshal is scheduled to inspect the Resource Center on Wednesday and we hope to be able to reopen the building later in the week.
  • Rapid progress on the temporary repairs to the Dental School continues to be made.
    • All 5 of the building buss ducts are now powered up.
    • The temporary switchgear has been delivered and the installation is approximately 50% complete.
    • Domestic water pumps are installed and operating, but the city water pressure is low and has been reported to the Sewerage &Water Board.
    • The temporary boiler is on site and the installation will begin shortly.